A square is enlarged by increasing the length of its sides by 10%. If the length of its sides was originally 6 cm, calculate the area of the enlarged square.
Answer by Theo(13342)

the original square is 6 by 6.
the area of the original square is 36 square units.
each side of the original square is increased by 10%.
this means each side is equal to 1.10 times what it was which means that each side of the enlarged square is equal to 1.10 * 6 = 6.6
the area of the enlarged square is equal to 6.6^2 = 43.56
if the side is multiplied by 1.10, then the area is multiplied by 1.10^2.
36 * 1.10^2 = 43.56.
this is because 1.10 * 6 * 1.10 * 6 is equal to 1.10 * 1.10 * 6 * 6 which is equal to 1.10^2 * 6^2.